# Differential Test Performance and Peer Effects

### v1

@@ 2.4 Estimation
Consistency and the asymptotic distribution of $\tilde{\theta}$ is discussed in Section 4 together with a discussion of the efficient estimator $\hat{\theta}$.
3 Empirical Results
@@
In this section, we apply our identification strategy to the Tennessee’s Project STAR (student-teacher achievement ratio) data, exploring peer effects among students in Kindergarten through Grade 3.999We downloaded the data from Harvard Dataverse (https://dataverse.harvard.edu/dataset.xhtml?persistentId=hdl:1902.1/10766), which differs slightly from the version used by Graham(2008) and Rose(2017). For example, they construct class ID using teacher characteristics while our data has class ID readily available. Their kindergarten sample has 6,172 students, and ours has 6,325. We believe our data set is more trustworthy as it matches the description in Word et al. (1990) in terms of sample size and summary statistics. Project STAR is a randomized experiment aiming at studying the impact of class size reduction on children’s development. The data set has been widely used in studies of class size, peer effects, teacher effectiveness and other education-related topics (e.g., Krueger 1999; Dee 2004; Graham 2008; Chetty et al. 2011; Mueller 2013). Here we briefly discuss the details which are relevant for our study. For a complete description of the project and data, see Word et al. (1990); Mosteller (1995); Boyd-Zaharias et al. (2007).
@@
Project STAR was carried out in Tennessee from 1985 to 1989, with additional data on the participants collected after the project ended. The experiments followed the kindergarten cohort of 1985 in participating schools for four years, from Kindergarten to the third grade. At the start of the 1985 academic year, participating schools randomly assigned kindergarten students and teachers into small classes (with an intended size of 13-17 students), regular classes (with an intended size of 22-25 students) and regular classes with a full-time teacher’s aide. In the actual implementation, small classes included 11 to 20 students, while regular classes (with or without aides) included 15 to 30 students. Randomness of the initial class assignment is ensured by the careful implementation of the random assignment protocol and has been examined by a number of studies, e.g., Krueger (1999); Chetty et al. (2011). However, nonrandom attrition, switching and migration might have happened in higher grades (Hanushek, 1999). Our analysis uses the sample of each of the four grades, as our estimator remains consistent even without random assignment. The bottom panel of Table 1 summarizes the number of students and classes in each grade. While there are 11601 students in the experiment in total, the sample size in each grade is between 6325 and 6840 due to students migrating in or out of the participating schools. The total count of schools is 79 at the Kindergarten level, but it decreases to 76 at the Grade 1 level and subsequently to 75 at the Grade 2-3 level. This decline can be attributed to schools withdrawing from the STAR program. There are between 325 and 340 classes in each grade, with 124 to 140 of them being small classes.
@@
We collect raw Stanford Achievement Tests (SAT) scores in mathematics (math), reading (read), listening (list) and word study skills (word) as test $t=1$ and test $t=2$ in our model. SAT is a nationally standardized test with scores comparable across grades. The test was administered on state specified examination days, occurring between late March and early April each year (Word et al., 1990; Krueger, 1999). The narrow time frame for the tests lends credence to the assumption that student’s ability and preparation remain stable over the testing window. SAT scores for these four subjects are available from Grade K to Grade 3. The mean and standard deviations of scores for the four SAT tests are in the third panel of Table 1. The four tests generally share comparable means, except that the mean listening score is much higher than others in Grade K. Notably, reading and word study skills scores have means that are close in all grades. With four test scores, there are 12 pairs of $y_{1}$ and $y_{2}$. We report results for only six unique combinations, as reversing the order of $y_{1}$ and $y_{2}$ does not change the estimate of $\rho$, transforms $f_{1}$ to its reciprocal, and will scale the coefficients of control variables by $-f_{1}$.
@@
Our main model includes a set of student and teacher characteristics, the mean and standard errors of which are summarized in the top panel of Table 1 by grade. Between $56.0\%$ and $59.2\%$ of the students receive a free lunch during the experimental period, an indicator of low family income. To account for the influence of race, we employ an indicator for black students, as more than 98% of the minority students are of black ethnicity. The share of black students ranges from $32.6\%$ to 34.7%. Girls make up $48.0\%$ to $48.6\%$ of the sample. We calculate age as of April 1st in each year, which is roughly the SAT examination date, using students’ dates of birth. Teacher’s characteristics are weighted by class size. Between 16.5% and 20.9% of the students are taught by black teachers. Meanwhile, 34.6% to 44.2% of the students have teachers with a master’s degree or higher. Average years of experience of the teachers for different grades are between 9.3 and 13.9.
@@
In our preferred specification, we include four types of control variables: school fixed effects (excluding one dummy for the first school), class type fixed effects (excluding the dummy for small classes), student characteristics (including indicators for free lunch, black ethnicity, female gender, and age) and teacher characteristics (including indicators for black teachers, having master’s degree or higher, and years of experience) and peer characteristics – specifically, the leave-out-mean of student characteristics. We start from the specification with only school fixed effects and subsequently introduce other types of control variables incrementally. We allow for heteroscedasticity of $u$ between small classes and regular classes (with or without teacher’s aide). Standard errors are clustered at the class level. The estimators for $f_{1}$ and $\rho$ are summarized in Tables 6 and 7 respectively.
@@
Before proceeding to discussing the main estimates of peer effects, we present several summary statistics in support of our model specifications and assumptions. Our estimator relies on the premise that tests $t=1$ and $t=2$ measure similar skills within a brief time frame. The combination of reading and word study skills aligns most closely with this narrative. To document the close connection between test scores, we use four measures to evaluate the correlation between various SAT scores from Grade K to Grade 3 in Table 2. The first two measures on the left panel are Spearman correlations between $y_{1}$ and $y_{2}$, and between $Q_{X}y_{1}$ and $Q_{X}y_{2}$. Here $Q_{X}=I-\underline{X}(\underline{X}^{\prime}\underline{X})^{-1}\text{$\underline{X}$}^{\prime}$ denotes the residual projection matrix of $\underline{X}$, where $\underline{X}$ encompasses the complete set of control variables: school fixed effects, class type fixed effect, student characteristics, teacher characteristics, and peer characteristics. The right panel of Table 2 reports the pseudo $R^{2}$ from 2SLS, i.e., the Spearman correlation between $y_{1}$ and $\hat{y}_{1}$, or between $y_{2}$ and $\hat{y}_{2}$. Here $\hat{y}_{1}=\tilde{f}_{1}y_{2}+\underline{X}\tilde{\delta}$, with $\tilde{f_{1}}$ and $\tilde{\delta}$ defined in (20) and (21). Similarly, $\hat{y}_{2}$ corresponds to the counterpart obtained by reversing the order of $y_{1}$ and $y_{2}$.  [^10] We also adjust for missing covariate values using a similar method as the one for the GMM estimator. That is replacing the peer averages of the observed individuals $\bar{x}_{c(-r)t}^{obs}=(n_{c}\bar{x}_{ct}-x_{crt}^{obs})/(n_{c}-1)$ with $\tilde{x}_{c(-r)t}^{obs}=(n_{c}\bar{x}_{ct}^{obs}-x_{crt}^{obs})/(n_{c}-1)$, where $\bar{x}_{ct}^{obs}$ is the average of all observed $x$.These adjustments make little difference to the results. We drop individuals with missing outcome variables from the sample. We discuss the procedure in more detail in Appendix E. The different correlation measures paint a consistent picture of the correlation between the six reported pairs as well as across grades. The listening score is generally a bit less correlated with other measures. The math scores are somewhat more correlated with reading and word skills than with listening. We find the strongest connection between word study skills and reading scores which exhibit substantial correlations of approximately 0.9 across all four measures and in all grades. This finding is consistent with these scores measuring skills that are both related to language comprehension. Correlations between other pairs range between 0.5 and 0.8. While a key identifying assumption, Assumption 2 is not directly testable because it relates to unobserved ability, finding strong correlation between test scores is consistent with a setting where test scores have common conditional means. If this interpretation of the results in Table 2 is correct then we would expect the word-read pair of scores to provide the best control for unobserved individual and class effects that may be confounding measures of peer effects at the classroom level. Our model implies that conditional on covariates correlation between tests only is due to unobserved interactive effects $\mu_{c}^{\ast}f_{t}.$ While not a test for correct specification, our results are consistent with $\mu_{c}^{\ast}f_{t}$ accounting for a large fraction of the variation in test scores.
Further evidence of the ability of test score differences to control for unobserved classroom and individual effects is the size of the estimated coefficient $f_{1}$. If scores measure closely related skills we expect $f_{1}$ to be close to one. In that scenario Assumption 2 is interpreted to imply that unobserved individual specific and classroom level effects captured by $\mu_{c}^{\ast}$ affect both scores in the same way. We report estimated values of $f_{1}$ for the set of test score pairs in Table 6. Point estimates for $f_{1}$ are obtained by our efficient GMM estimator for all grades and for specifications with just a school effect in Column (1), school and class type effects in Column (2), with a full set of controls excluding contextual peer effects in Column (3) and the full set of controls in Column (4). The coefficient $f_{1}$ is estimated with high precision in Columns (1) through (3) and somewhat less precisely in Column (4).The point estimates are generally close to one with a few exceptions, such as for the listening scores in grade K and specifications (1)-(3) as well as the listening scores in grades two and three for specification (4). The parameter estimates for the read-word score combination are remarkably stable across all specifications and across all grades, are close to one and estimated precisely. This is further evidence that the read-word score pair is best suited to control for unobserved class level effects.
@@
We explore one additional approach to evaluate the plausibility of Assumption 2. If unobserved individual and class effects $\mu_{c}^{\ast}$ are stable over time then scores from previous years should be good predictors for current year scores. Here, it is important to note that our theory applies to pairs of closely related tests, and that $\mu_{c}^{\ast}$ is not assumed to be the same for different pairs. Thus our theory does not assume $\mu_{c}^{\ast}$ to be time invariant between grades. Conceptually, $\zeta^{*}$ measures unobserved ability just prior to taking the tests in the current year. As a result, $\mu_{c}^{\ast}$ similarly reflects only current year unobserved ability and other classroom level effects coming from teachers and peers. Unobserved effects $\mu_{c}^{\ast}$ are expected to vary year by year for various reasons, including the fact that instruction during the year improves skills, that teachers may change and that students may move to different classes. However, it is expected that at least some components of $\mu_{c}^{\ast}$ that are more closely related to individual student ability are time invariant. We examine this hypothesis by measuring the explanatory power of lagged scores on current scores. The regression sample includes stayers in each grade, defined as individuals participating in the STAR project for both the previous and the current grades to ensure availability of both current and lagged scores. The Grade 1 sample includes 66% stayers, while Grade 2 and 3 samples have 73.8% and 79.6% respectively. Among the stayers in Grade 1, 88.7% stay in the same type of classes (small v.s. regular(w/wo) aide) in the same schools. This proportion exceeds 93% for Grade 2 and Grade 3.111111Based on the authors’ calculation using the STAR data. The share is lower if we consider regular and regular/aid as different types, due to a random reallocation of students in regular and regular/aid classes to these two types within school in Grade 1. However, the literature generally finds no significant impact of teacher’s aide (e.g., Krueger, 1999). The stability of school and class types among the large share of stayers fosters a steady learning environment. While newcomers and leavers might marginally influence peer quality, the allocation of newcomers to different classrooms is random within schools. Consequently, when accounting for school fixed effects, we anticipate a consistent level of peer quality over time.
@@
The relatively stable environment that these students are in suggests that $\mu_{c}^{\ast}$ and $f_{t}$ should contain significant time invariant components that in turn imply high correlation between observed scores and lagged scores. As a result we expect to see relatively high $R^{2}$ in predictive regressions with lagged scores as controls in addition to the observed characteristics. Consequently, we regress test outcomes on the full set of control variables (school and class type fixed effects, students, teacher and peer characteristics and a constant term), and then add lagged outcomes of the same or different type and their peer averages to assess changes in $R^{2}$. The results are summarized in Table 3. The results show the predictive power of lagged scores for the same test in bold, as well as when using other tests instead, for grades one through three. The predictive power of lagged scores is documented by comparing the $R^{2}$ with and without adding lags. For example, for the word score, adding its own lag improves the $R^{2}$ from .244 to .435 in grade one, from .25 to .539 in grade two and from .229 to .488 in grade three. Using the lagged reading score, which is the score of the test most closely related to the word score, leads to increases in the $R^{2}$ that are even slightly higher. Overall, the coefficients on lagged scores when using the own lag of word as well as the lagged reading score are large in magnitude and highly statistically significant. These findings hold up for the other test score pairs with similar magnitudes and statistical significance. As may be expected, the predictive effects are slightly weaker for listening scores than for math, read and word. The results are further evidence that the data distributions are consistent with key features of our model.
Our analysis thus far has documented that pairs of tests scores are highly correlated both contemporaneously as well as over the course of one school year. This evidence is consistent with restrictions imposed by Assumption 2 and the hypothesis that closely related scores depend on the same unobserved components. A caveat is that closely related scores may remove too much common variation when quasi-differences of these scores are used and thus render statistical evidence about peer effects imprecise. Table 4 reports pseudo $R^{2}$ measures of model estimates for the six pairs of scores and model specifications (1) through (4) with column (1) only controlling for school fixed effects, and column (4) being the full specification with school and class type effects, student and teacher characteristics as well as peer characteristics.
The observed pseudo $R^{2}$ measures are generally below 0.2 with a few exceptions for Grade $K$. We also see that the $R^{2}$ is a bit higher for models with listening scores. Moreover and as expected the pseudo $R^{2}$ generally increases with larger models, or in other words by moving from Column (1) to (4). The read-word score pair often has the lowest $R^{2}$ which is an indication that the differential score in these specifications already controls for most of the variation explained by the covariates. We take this as further evidence that read-word best fits the model assumptions we impose.
Table 5 presents standard deviations of $Q_{x}y$ for four test scores and $Q_{x}(y_{1}-f_{1}y_{2})$ for six pairs of test scores for Grade K to 3. The left and right panels are for small and regular classes respectively. Control variables for specifications (1)-(4) are the same as those in Table 4, as discussed above. Estimates of $f_{1}$ are obtained from our efficient GMM estimator and are presented in Table 6. We see that in Grade K and Grade 1, the ratio of standard deviations of $Q_{x}(y_{1}-f_{1}y_{2})$ to that of $Q_{x}y$ is around 1/2 for most pairs of tests and 1/5 to 1/4 for the read-word pair. In Grade 2 and 3, this ratio is around 1/3 for most pairs and about 1/5 for the read-word pair. This shows that quasi-differencing removes a significant part of the variation in the test scores, especially for the read-word pair, consistent with the term $\mu^{*}f_{1}$ playing a major role in determining test performance. At the same time, the results in Table 4 indicate that the inclusion of controls does not eliminate much of the remaining variation not captured by fixed effects. The conclusion is that there is between 20 to 30 percent of test score variation remaining for the identification of $\rho$.

### v2

@@ 3.3 Estimation
We prove identification in Section 5. Consistency and the asymptotic distribution of $\tilde{\theta}=\left(\tilde{\rho},\tilde{f}_{1},\tilde{\delta}^{\prime}\right)^{\prime}$are discussed in Sections C and D in the online appendix together with a discussion of the efficient estimator $\hat{\theta}$ defined below.
The estimator $\tilde{\theta}$ is inefficient as it is constructed from the residuals $u_{1}-f_{1}u_{2}$ in (12), which are not homoscedastic. To develop efficient estimators, denote the variance-covariance matrix of $u_{1}-f_{1}u_{2}$ as $\Omega(\gamma)$, where $\gamma$ are parameters to be specified later. Theorem 2.1 indicates that $\Omega(\gamma)$ is a diagonal matrix. The form of $\Omega(\gamma)$ and $\gamma$ are discussed under Assumption 5 as well as Assumption 9 in the online appendix.
Premultiplying (12) by $V\left(\rho,\gamma\right)=\Omega(\gamma)^{-1/2}(I+\rho M)^{-1}$ then yields the following equation
$y^{+}(\theta,\gamma)=X^{+}(\theta,\gamma)\delta+u^{+}$ (17)
where $y^{+}(\theta,\gamma)=V\left(\rho,\gamma\right)(y_{1}-f_{1}y_{2}),$ $X^{+}(\theta,\gamma)=V\left(\rho,\gamma\right)\underline{X},$ and $u^{+}=\Omega(\gamma)^{-1/2}(u_{1}-f_{1}u_{2})$ and $\operatorname{Var}\left(u^{+}|\mathcal{Z}_{n}^{*}\right)=I_{n}$. In light of (17) we define
[…]
and observe that $u^{+}(\theta,\gamma)=u^{+}$ at the true parameters. The transformation leading to $u^{+}$ consists of three components. Quasi-differencing eliminates classroom and individual effects. The operator $(I+\rho M)^{-1}$ removes cross-sectional correlation due to peer effects and can be understood as a form of spatial Cochrane-Orcutt transformation. Finally, the operator $\Omega(\gamma)^{-1/2}$ scales the spatially uncorrelated residuals to unit variance in the case of heteroskedasticity.
We next provide an outline of our GMM estimation methodology. For clarity we explicitly denote in the following the true parameter vectors as $\theta_{0}=(\rho_{0},f_{1,0},\delta_{0}^{\prime})^{\prime}$ and $\gamma_{0}$. Our GMM estimator [^5] Rigorous consistency and asymptotic normality results as well as additional assumptions needed to establish these properties are given in Sections C and D in the online appendix. is based on both linear and quadratic moment conditions.
$m_{n}(\theta,\gamma)=\left(\begin{array}[]{c}m_{n}^{l}(\theta,\gamma)\\ m_{n}^{q}(\theta,\gamma)\end{array}\right)=n^{-1/2}\left(\begin{array}[]{c}H^{\prime}u^{+}(\theta,\gamma)\\ u^{+\prime}(\theta,\gamma)Au^{+}(\theta,\gamma)\end{array}\right),$ (19)
where the functions $m_{n}^{l}(\theta,\gamma)$ and $m_{n}^{q}(\theta,\gamma)$ denote the linear and quadratic moment conditions respectively. Under the maintained assumptions of this paper the results in Kuersteiner and Prucha (2020), Theorem 1, imply that the linear moment function is uncorrelated with the quadratic moments. We have
[…]
$Q_{n}\left(\theta,\gamma\right)$ $=\frac{1}{n}u^{+}(\theta,\gamma)^{\prime}H\left(H^{\prime}H\right)^{-1}H^{\prime}u^{+}(\theta,\gamma)+\frac{1}{n}\frac{\left(u^{+}(\theta,\gamma)^{\prime}Au^{+}(\theta,\gamma)\right)^{2}}{2\operatorname{tr}\left(A^{2}\right)}$ (23)
such that the GMM estimator for $\theta_{0}$ is defined as $\hat{\theta}\left(\gamma\right)=\operatorname{argmin}_{\theta}Q_{n}\left(\theta,\gamma\right)$ for a fixed value of $\gamma.$ In applications $\gamma$ is replaced with a first step consistent estimator.
4 Peer Effects in Project STAR
4.1 Data
@@
In this section, we apply our identification strategy to the Tennessee’s Project STAR (student-teacher achievement ratio) data, exploring peer effects among students in Kindergarten through Grade 3.777We use STAR data obtained from Harvard Dataverse (https://dataverse.harvard.edu/dataset.xhtml?persistentId=hdl:1902.1/10766), which differs slightly from the version used by Graham(2008) and Rose(2017). For example, they construct class ID using teacher characteristics while our data has class ID readily available. Their kindergarten sample has 6,172 students, and ours has 6,325. We prefer our data as it matches the description in Word et al. (1990) in terms of sample size and summary statistics. Project STAR is a randomized experiment aiming at studying the impact of class size reduction on children’s development. The data set has been widely used in studies of class size, peer effects, teacher effectiveness and other education-related topics (e.g., Krueger 1999; Dee 2004; Graham 2008; Chetty et al. 2011; Mueller 2013). Here we briefly discuss the details which are relevant for our study. For a complete description of the project and data, see Word et al. (1990); Mosteller (1995); Boyd-Zaharias et al. (2007).
@@
Project STAR was carried out in Tennessee from 1985 to 1989, with additional data on the participants collected after the project ended. The experiments followed the kindergarten cohort of 1985 in participating schools for four years, from Kindergarten to the third grade. At the start of the 1985 academic year, participating schools randomly assigned kindergarten students and teachers into small classes , regular classes and regular classes with a full-time teacher’s aide. In the actual implementation, small classes included 11 to 20 students, while regular classes (with or without aides) included 15 to 30 students. Randomness of the initial class assignment is ensured by the careful implementation of the random assignment protocol and has been examined by a number of studies, e.g., Krueger (1999); Chetty et al. (2011). However, nonrandom attrition, switching and migration might have happened in higher grades (Hanushek, 1999). Our analysis uses the sample of each of the four grades, as our estimator remains consistent even without random assignment. [^8] A Table with summary statistics for each grade can be found in the online appendix. While there are 11601 students in the experiment in total, the sample size in each grade is between 6325 and 6840 due to students migrating in or out of the participating schools. The total count of schools is 79 at the Kindergarten level, but it decreases to 76 at the Grade 1 level and subsequently to 75 at the Grade 2-3 level. This decline can be attributed to schools withdrawing from the STAR program. There are between 325 and 340 classes in each grade, with 124 to 140 of them being small classes.
@@
Project STAR data contains tests scores for mathematics, reading, listening and word study skills. We use raw Stanford Achievement Tests (SAT) scores in word study skills (word) and reading (read) as test $t=1$ and test $t=2$ respectively in our model, as they measure closely related skills and hence are more likely to satisfy our assumptions. This choice is supported by diagnostic checks we report in Tables 1 2 and Table 6 in the appendix. Reversing the order of $y_{1}$ and $y_{2}$ does not change the estimate of $\rho$. Results using scores in mathematics (math) and listening (list) are discussed in the online appendix as a comparison. SAT is a nationally standardized test with scores comparable across grades. The test was administered on state specified examination days, occurring between late March and early April each year (Word et al., 1990; Krueger, 1999). The narrow time frame for the tests lends credence to the assumption that student’s ability and preparation remain stable over the testing window. The reading and word study skills scores have means that are close in all grades.
@@
Our preferred model includes a set of student and teacher characteristics. Between $56.0\%$ and $59.2\%$ of the students receive a free lunch during the experimental period, an indicator of low family income. To account for the influence of race, we employ an indicator for black students, as more than 98% of the minority students are of black ethnicity. The share of black students ranges from $32.6\%$ to 34.7%. Girls make up $48.0\%$ to $48.6\%$ of the sample. We calculate age as of April 1st in each year, which is roughly the SAT examination date, using students’ dates of birth. Teacher’s characteristics are weighted by class size. Between 16.5% and 20.9% of the students are taught by black teachers. Meanwhile, 34.6% to 44.2% of the students have teachers with a master’s degree or higher. Average years of experience of the teachers for different grades are between 9.3 and 13.9.
@@
In our preferred specification, we include four types of control variables: school fixed effects (excluding one dummy for the first school), class type fixed effects (excluding the dummy for small classes), student characteristics (including indicators for free lunch, black ethnicity, female gender, and age) and teacher characteristics (including indicators for black teachers, having master’s degree or higher, and years of experience) and peer characteristics – specifically, the leave-out-mean of student characteristics. We start from the specification with only school fixed effects and subsequently introduce other types of control variables incrementally. We find that our results are not very sensitive to the inclusion of additional controls. This is reassuring in that our differencing strategy of closely related scores removes most of the systematic variation in scores that is related to observed and unobserved individual and classroom level factors affecting outcomes. We allow for heteroscedasticity of $u$ between small classes and regular classes (with or without teacher’s aide). Standard errors are clustered at the class level.
4.2 Plausibility of Assumptions
@@
Before proceeding to discussing the main estimates of peer effects, we present several summary statistics in support of our model specifications and assumptions. Our estimator relies on the premise that word score $y_{1}$ and reading score $y_{2}$ measure similar skills within a brief time frame. To document the close connection between test scores, we use four measures to evaluate the correlation between $y_{1}$ and $y_{2}$ from Grade K to Grade 3 in Table 1. The first two measures on the left panel are Spearman correlations between $y_{1}$ and $y_{2}$, and between $Q_{X}y_{1}$ and $Q_{X}y_{2}$. Here $Q_{X}=I-\underline{X}(\underline{X}^{\prime}\underline{X})^{-1}\text{$\underline{X}$}^{\prime}$ denotes the residual projection matrix of X, where X encompasses the complete set of control variables: school fixed effects, class type fixed effect, student characteristics, teacher characteristics, and peer characteristics. The right panel of Table 1 reports the pseudo $R^{2}$ from 2SLS, i.e., the Spearman correlation between $y_{1}$ and $\hat{y}_{1}$, or between $y_{2}$ and $\hat{y}_{2}$. Here $\hat{y}_{1}=\tilde{f}_{1}y_{2}+\underline{X}\tilde{\delta}$, with $\tilde{f_{1}}$ and $\tilde{\delta}$ defined in (14) and (15) being the 2SLS estimator when $y_{2}$ is instrumented by 1. Similarly, $\hat{y}_{2}$ corresponds to the counterpart obtained by reversing the order of $y_{1}$ and $y_{2}$. [^9] We also adjust for missing covariate values using a similar method as the one for the GMM estimator. That is replacing the peer averages of the observed individuals $\bar{x}_{c(-r)t}^{obs}=(n_{c}\bar{x}_{ct}-x_{crt}^{obs})/(n_{c}-1)$ with $\tilde{x}_{c(-r)t}^{obs}=(n_{c}\bar{x}_{ct}^{obs}-x_{crt}^{obs})/(n_{c}-1)$, where $\bar{x}_{ct}^{obs}$ is the average of all observed $x$.These adjustments make little difference to the results. We drop individuals with missing outcome variables from the sample. We discuss the procedure in more detail in Section G of the online appendix. We find that the word study skills and reading scores exhibit substantial correlations of approximately 0.9 across all four measures and in all grades. This finding is consistent with these scores measuring skills that are both related to language comprehension. On the other hand, correlations between other pairs are smaller and range between 0.5 and 0.8, as shown in Table 6 in the Appendix. While a key identifying assumption, Assumption 2 is not directly testable because it relates to unobserved ability, finding strong correlation between test scores is consistent with a setting where test scores have common unobserved individual and class effects that may be confounding measures of peer effects at the classroom level. Our model implies that conditional on covariates correlation between tests only is due to unobserved fixed effects $\mu_{c}^{\ast}$ and peer effects working through idiosyncratic errors $u_{it}.$ While not a test for correct specification, our results are consistent with $\mu_{c}^{\ast}$ accounting for a large fraction of the variation in test scores.
@@
We next explore an additional approach [^10] Two more specification checks are reported in the online appendix. to evaluate the plausibility of Assumption 2. If unobserved individual and class effects $\mu_{c}^{\ast}$ are stable over time then scores from previous years should be good predictors for current year scores. Here, it is important to note that our theory applies to pairs of closely related tests, and that $\mu_{c}^{\ast}$ is not assumed to be the same for different pairs. Thus our theory does not assume $\mu_{c}^{\ast}$ to be time invariant between grades. Conceptually, $\zeta^{*}$ measures unobserved ability just prior to taking the tests in the current year. As a result, $\mu_{c}^{\ast}$ similarly reflects only current year unobserved ability and other classroom level effects coming from teachers and peers. Unobserved effects $\mu_{c}^{\ast}$ are expected to vary year by year for various reasons, including the fact that instruction during the year improves skills, that teachers may change and that students may move to different classes. However, it is expected that at least some components of $\mu_{c}^{\ast}$ that are more closely related to individual student ability are time invariant. We examine this hypothesis by measuring the explanatory power of lagged scores on current scores. The regression sample includes stayers in each grade, defined as individuals participating in the STAR project for both the previous and the current grades to ensure availability of both current and lagged scores. The Grade 1 sample includes 66% stayers, while Grade 2 and 3 samples have 73.8% and 79.6% respectively. Among the stayers in Grade 1, 88.7% stay in the same type of classes (small v.s. regular(w/wo) aide) in the same schools. This proportion exceeds 93% for Grade 2 and Grade 3.111111Based on the authors’ calculation using the STAR data. The share is lower if we consider regular and regular/aid as different types, due to a random reallocation of students in regular and regular/aid classes to these two types within school in Grade 1. However, the literature generally finds no significant impact of teacher’s aide (e.g., Krueger, 1999). The stability of school and class types among the large share of stayers fosters a steady learning environment. While newcomers and leavers might marginally influence peer quality, the allocation of newcomers to different classrooms is random within schools. Consequently, when accounting for school fixed effects, we anticipate a consistent level of peer quality over time.
@@
The relatively stable environment that these students are in suggests that $\mu_{c}^{\ast}$ should contain significant time invariant components that in turn imply high correlation between observed scores and lagged scores. As a result we expect to see relatively high $R^{2}$ in predictive regressions with lagged scores as controls in addition to the observed characteristics. Consequently, we regress test outcomes on the full set of control variables (school and class type fixed effects, students, teacher and peer characteristics and a constant term), and then add lagged outcomes of the same or different type and their peer averages to assess changes in $R^{2}$. The results are summarized in Table 2. The results show the predictive power of lagged scores for the same test in bold, as well as when using other tests instead, for grades one through three. The predictive power of lagged scores is documented by comparing the $R^{2}$ with and without adding lags. For example, for the word score, adding its own lag improves the $R^{2}$ from 0.244 to 0.435 in grade one, from 0.25 to 0.539 in grade two and from 0.229 to 0.488 in grade three. Using the lagged reading score leads to increases in the $R^{2}$ that are even slightly higher. Overall, the coefficients on lagged scores when using the own lag of word as well as the lagged reading score are large in magnitude and highly statistically significant. The results are further evidence that the data distributions are consistent with key features of our model.
Table 3 presents estimates and standard errors (in parentheses) for $f_{1}$ (left panel) and the peer effects parameter $\rho$ (right panel).  [^12] Results for contextual peer effects are reported in Table 10 in the online appendix. Estimates for other pairs of scores are in Table 7 in the online appendix. These estimates are obtained by our efficient GMM estimator for all grades and for specifications with just a school effect in Column (1), school and class type effects in Column (2), with a full set of controls excluding contextual peer effects in Column (3) and the full set of controls in Column (4).
The point estimates of $f_{1}$ are remarkably stable across all specifications and across all grades. The size of the estimated coefficient $f_{1}$ is close to one and estimated precisely. This is further evidence that word and reading scores measure closely related skills and that the test score differences control unobserved classroom and individual effects. Assumption 2 postulates that unobserved individual specific and classroom level effects captured by $\mu_{c}^{\ast}$ affect both scores in the same way. This is credible when both scores measure essentially the same skills. Our empirical finding of $f_{1}$ close to one for the word-read pair is consistent with this interpretation of the data. Other score pairings reported in the appendix have estimated $f_{1}$ parameters that are less reliably close to one, lending further support for our choice of focusing on word-read as our preferred specification.
4.3 Peer Effects